pls. help me solve this prob. tnx.
prob.1. Mary and Nancy set off on a 220- km jeepney. Nancy started the journey 15minutes earlier than mary. They reached the destination at the same time. If Nancy's speed was 80km/hr, find the Mary's speed.

Answer by mananth(16949)

let Mary's speed = x kph
Nancy's speed = 80 kph
They reached at the same time
220/80 -220/x =1/4
LCD = 80x
(220x-220*80)/80x=1/4
220x-17600=80x/4
(220x-17600)=20x
220x-20x=17600
200x= 17600
x=17600/200
x=88 kph Mary's speed

m.ananth@hotmail.ca